OpenStudy (anonymous):

solve the equation cos(2x)= √2-cos(2x)

ganeshie8 (ganeshie8):

square both sides, plugin cos(2x) = u, solve the quadratic

OpenStudy (anonymous):

You add cos(2x) to both sides: \[\cos(2x)+\cos(2x) = \sqrt{2}\] \[2\cos(2x) = \sqrt{2}\] \[\cos(2x) = \frac{\sqrt{2}}{2}\] \[\cos(2x)=cos( \frac{π}{4})\] \[2x=2kπ+\frac{π}{4} (Equation\ 1)\] or \[2x=2kπ-\frac{π}{4}(Equation\ 2)\] And you solve both for x.

OpenStudy (anonymous):

after pluggin in cos2x=u,u get a quadratic equation in u, u^2+u-2=0,solving it,u get u=2,-1.therefore cos2x=2,-1.but cos2x,with its range restricted to [-1,1] cannot be equal to 2.so,cos2x=-1.which gives